OpenStudy (anonymous):

Find the center radius form of the equation of the circle described and graph it. center (0,-1) radius 6

OpenStudy (anonymous):

how do you set it up? I dont get it

OpenStudy (anonymous):

i dont know the formula

OpenStudy (anonymous):

Look at the other post: http://openstudy.com/study#/updates/505a6c00e4b03290a415346e

OpenStudy (anonymous):

(x-h)^2 + (y-k)^2 = r^2

OpenStudy (anonymous):

center ( h, k) = (0,-1)

OpenStudy (anonymous):

(x-0)^2+(y-1)^2=r^2

OpenStudy (anonymous):

x^2-0)^2+(y-1)^2=6^2

OpenStudy (anonymous):

x² + ( y+1)² = 36 That's how you plug the numbers into the circle formula!

OpenStudy (anonymous):

thank you